IONISING RADIATION QUESTION 2

The risk for stochastic effects when a person receives a non-whole body dose can be described by:

Select one:

a. The activity.

b. The absorbed dose.

c. The effective dose.

d. The equivalent dose.

The effective dose from an x-ray point source is measured to be D mSv. If the distance to the source is doubled and the exposure time is halved, what is the new effective dose?

Select one:

a. D mSv

b. D/2 mSv

c. D/4 mSv

d. D/8 mSv

Lead sheet is commonly used as a shielding material for x-rays. When a 1mm thick piece of lead is placed in front of an x-ray tube, it is observed that the fraction of x-rays transmitted is 25%. What total thickness is needed to reduce the fraction transmitted to approximately 1%?

Select one:

a. 4.4 mm

b. 3.3 mm

c. 2.4 mm

d. 2.9 mm
Answer #1

Stochastic effects are the effects of whose probability of occurence increases with dose. Effective dose is a tissue weighted sum of the equivalent doses  in all specified tissues and organs of the human body and represents the stochastic health risk to the whole body.

Answer : Effective Dose

2. Effective dose =kt/d2 =D

Where k is constant, t is exposure time and d is distance

We have , t =2t0, d=d0/2

Hence effective dose=D/2 mSv

3. - ua Attunation of x-rays is given by I= I e where,I transmitted intensity T- Original intensity u= constant thickness 2 - Foe-u 4 e u 4 In eu In 4 u= 1.386 Now, let ned, I = I. 100 . - I. 100 e-ud -ud e 100 100 Eetud In see in 100 end In ud ud 4.605